Please use all 5 places of decimal when working with probabilities and please use at least 3 significant figures when dealing with proportions. For larger numbers, please use two places of decimal. 1. Professor Marc has chosen a random sample of 10 CSI students in order to test his hypothesis that the average student will pay $40 for a textbook. His sample shows an average amount these students are willing to pay for the textbook of $37 with a standard deviation of about $12. Using a t-test and α-0.01, test the hypothesis that the average amount students at CSI are willing to pay for a textbook is $40.
Answer #1

here sample size=n=10, we want to test the

null hypothesis H0:mu=40 and

alternate hypothesis Ha:μメ40 ( this is two tailed test)

the statistic t=|(ar{x}-mu)|/(s/sqrt(n))=|(37-40)|/(12/sqrt(10))=0.79 with n-1=10-1=9 df

the critical t(0.01,9)=3.25 is greater than calculated t=0.79, so we fail to reject H0 and conclude that students are willing to pay for a textbook is $40.